jhonyy9 (jhonyy9):

- so the collatz's conjecture said ,,... if is even divide by 2 and if is odd multiplie by 3 and add 1 .... so to the end allways will result 1. example. 1*3 +1 = 3+1=4 what is power of 2 so what mean that dividing by 2 =2/2 =1 so we get 1 2 is even so 2/2 =1 3 odd so 3*3 +1 = 10 even so 10/2=5 what is odd so 5*3 +1 = 16 what is power of 2 what mean that on the end will result 1 4=2^2 so will result 1 5 odd so 5*3+1=16 = 2^4 power of 2 so will result 1 -so than we need proving that in case of every odds will result power of 2 using this method prove that allways will be a natural number n for what (2n+1)*3 +1 will result power of 2

OpenStudy (primeralph):

(2n+1)*3 +1 A statement ........ equals what?

jhonyy9 (jhonyy9):

yes i like continue it here every odds we can writing in the form of 2n+1 yes ? so (2n+1)*3+1=6n+3+1=6n+4 =2(3n+2) so what mean that in case of every odds we will get like result even

jhonyy9 (jhonyy9):

so for we can proving the collatz"s conjecture we need proving that using this method of dividing by 2 when is even and multiplying by 3 and adding 1 when is odd so in this case on the end will result allways power of 2 (2n+1)*3 +1 = 6n +3 +1 = 6n +4 = 2(3n +2) /2 = 3n +2 so for we can proving what we have wrote above us reductio add absurdum method suppose that 3n+2 not equale never power of 2 but for n=2 we get 3*2 +2 = 8 = 2^3 Q:E:D: opinions please ! thank you very much

jhonyy9 (jhonyy9):

@saifoo.khan @satellite73 @dumbcow @ganeshie8

jhonyy9 (jhonyy9):

@ash2326 @sauravshakya

jhonyy9 (jhonyy9):

@Hero

OpenStudy (anonymous):

well you have just proved following statement wrong: for all n , (3n+2) can never be written as power of 2. but this does not prove collatz conjecture

OpenStudy (anonymous):

Actually once I also tried to think on this problem....... approached from various ways but got nothing....... I guess if u prove after some steps u will always reach a smaller number than a number that u started with then it will prove the conjecture

jhonyy9 (jhonyy9):

ty but if you check it for an standard form of odd numbers why you not can using the reductio add absurdum method of ,... and why not can being this accepted ?

OpenStudy (anonymous):

what do you mean by reductio add absurdum method?

jhonyy9 (jhonyy9):

what have wrote there because (2n+1)*3+1 will result allways even what divide by 2 we get 3n+2 so this 3n+2 suppose not will be never power of 2 but for n=2 we get 8 = 2^3

jhonyy9 (jhonyy9):

@Hero @ash2326 @radar @dumbcow opinion

jhonyy9 (jhonyy9):

@amistre64

jhonyy9 (jhonyy9):

@cwrw238 opinion

jhonyy9 (jhonyy9):

@jim_thompson5910

OpenStudy (radar):

Sorry my opinion does not count.

jhonyy9 (jhonyy9):

but do you understand this proof by reductio add absurdum ?

jhonyy9 (jhonyy9):

and how is possibille that ,,persons with title of champion dont know the reductio ad absurdum proof method ?

OpenStudy (anonymous):

it's typically called proof by contradiction

jhonyy9 (jhonyy9):

@oldrin.bataku yes i think that you see it right but when i have learned this was the name of this proof method -- reductio ad absurdum -- yes ,what mean proof by contradiction ty ...and what is your opinion from this proof by contradiction ? you see it correct or ...?... your opinion ?

OpenStudy (anonymous):

I think you are misinterpreting the Collatz conjecture

jhonyy9 (jhonyy9):

how do you think this and why ?

jhonyy9 (jhonyy9):

Collatz said

jhonyy9 (jhonyy9):

take indifferent number : than even divide it by 2 than odd multiplie it by 3 and add 1 so the final result will be allways 1 yes ?

jhonyy9 (jhonyy9):

@radar you see it the same too ? please what is your opinion from ?

OpenStudy (radar):

Yes, for example if the number was 9, the sequences would be 9, 28, 14, 7, 22, 11, 34, 17, 52, 26, 13, 40, 20, 10, 5, 16, 8, 4, 2, 1 at least I know that one does, but I don't know if there may be a number where it won't go that way.

OpenStudy (radar):

Like a real large number, as they have tested numbers way way way up there, and so far the sequence has gone to 1.

OpenStudy (radar):

I would assume a negative number would terminate at -1, still a 1 (I think)

OpenStudy (radar):

Gotta run, my meat loaf is ready.

OpenStudy (cwrw238):

some numbers take a lot of steps to terminate e g n=27 takes 111 steps before reaching 1 others , a lot longer like the Riemman hypothesis this is one of the conjectures that hasn't been proved yet

OpenStudy (cwrw238):

proof by contradiction is powerful technique you start by assuming the opposite of the hypothesis is true eg to prove that sqrt 2 is irrational assume that it can be written as a/b (that its rational) then by using logical steps you come to a statement that cannot be true (thats absurd - hence the 'redact ad absurbium.') thus you conclude that your assumption cannot be correct therefore the original hypothesis must be true if you look up this proof you'll see the details in the case of collasz conjecture you'd start by assuming the process can result in a number other than 2

OpenStudy (cwrw238):

* other than 1

jhonyy9 (jhonyy9):

ty so than like finaly result you consider it proven by reductio ad absurdum method ?

OpenStudy (cwrw238):

i dont know - i'll have to look at it later no time now

OpenStudy (anonymous):

if you assume 3n+2 is not a power of 2 for all natural numbers and use contradiction method, then you are just proving there exist at least one integer n for which 3n+2 is a power of 2

OpenStudy (cwrw238):

right